OpenStudy (anonymous):

Gia’s gross annual salary is $46,392. What is the recommended maximum monthly rent she should pay?

OpenStudy (texaschic101):

maximum housing expense ratio = (annual salary x 0.28) / 12 maximum housing expense ratio = (46,392 x 0.28) / 12 = 12989.76/12 = 1082.48 I hope thats right.....you might better get a second opinion Front-end ratio: The housing expense, or front-end, ratio shows how much of your gross (pretax) monthly income would go toward the mortgage payment. As a general guideline, your monthly mortgage payment, including principal, interest, real estate taxes and homeowners insurance, should not exceed 28 percent of your gross monthly income. To calculate your housing expense ratio, multiply your annual salary by 0.28, then divide by 12 (months). The answer is your maximum housing expense ratio.